# Spatial Reasoning with Vision-Language Models in Ego-Centric Multi-View Scenes

### v1

@@ 5 Evaluation on Ego3D-Bench
In this section, we present a comprehensive evaluation of VLMs on Ego3D-Bench. We organize our experiments into four parts (1) benchmarking generalist VLMs (i.e., models trained for general vision-language tasks), (2) benchmarking 3D-VLMs (i.e., models trained specifically for 3D spatial understanding), (3) benchmarking VLM+Depth+REC (i.e., generalist VLMs augmented with depth and REC tools), and (4) ablation studies. We use a fixed R1-style prompt Guo et al. (2025) in the evaluations of all models (details in the Appendix). We use Grounding-Dino-Base Research (2023) as the REC model and Depth-Anything-V2-Metric-Large Contributors (2024) as the metric depth estimator in all experiments.
5.1 Benchmarking Generalist VLMs
From the closed-source models, we use GPT-4o OpenAI (2024) and Gemini-2-Flash DeepMind (2023). From the open-source models, we select three competitive families of open-source models: InternVL3 et al. (2025), Qwen2.5-VL Qwen et al. (2025), and Ovis2 Lu et al. (2024). Qualitative examples, inference time analysis, and numerical results on more open-source models are given in the Appendix.
Table 1: Comparison results of generalist VLMs and our method (Ego3D-VLM) on Ego3D-Bench.
@@ 5.1 Benchmarking Generalist VLMs
Performance Analysis. Figure 6 illustrates the average performance of the four leading models (GPT-4o, Gemini-1.5-Pro, InternVL3-78B, and Qwen2.5-VL-72B) with and without the integration of Ego3D-VLM across all multiple-choice questions. Travel time, localization, and object-centric absolute distance are the most challenging categories for VLMs with 40%-45% average accuracy with the largest models of distinct families. We argue that these categories require intricate spatial reasoning and the ability to construct an accurate mental map based on the relative positioning of objects. While humans can effortlessly build such mental maps and achieve perfect accuracy, current VLMs struggle to replicate this level of spatial understanding, highlighting a key area for further development.
Figure 6: Average performance of leading VLMs (GPT-4o, Gemini-1.5-Pro, InternVL3-78B, and Qwen2.5-VL-72B) with/without Ego3D-VLM vs. chance and human levels on each category of Ego3D-Bench.
In the absolute distance measurement task, the incorporation of 3D location data through the cognitive map substantially narrows the gap between model performance and human level accuracy. Notably, for object-centric absolute distance estimation, VLMs augmented with Ego3D-VLM even surpass human performance. This outcome is anticipated, as human approximation of 3D distances in object-centric scenarios is prone to substantial error without explicit 3D spatial information. Conversely, in the localization task, models continue to fall short of human proficiency, even with cognitive map support.
Blind VLM Performance. This baseline evaluates how much spatial reasoning can be achieved by VLMs using only textual input, without any visual information, relying solely on their world knowledge Majumdar et al. (2024). We report the average performance of GPT-4o and Gemini-1.5-Pro models. The blind VLMs perform 5% worse than vision-enabled VLMs (53.8% vs. 58.8%) and 16.4% better than chance level. These results are consistent with findings from prior single-view and video-based spatial reasoning benchmarks Cheng et al. (2024); Yang et al. (2024).
Table 2: Performance of cognitive map fed to LLM.
@@
Blind Ego3D-VLM Performance. In this section, we perform an experiment by feeding the cognitive map to a text only LLM. Since InternVL3-8B/14B use Qwen2.5-7B/14B as their backbone LLM, we use the same model as the text only LLM for fair comparison. Table 2 shows that Ego3D-VLM with VLM outperforms Ego3D-VLM with LLM since VLMs can ignore false positives in the cognitive map and remain robust when false negatives occur, whereas LLMs suffer performance drops in those cases.
@@ 5.2 Benchmarking 3D-VLMs
3D-VLM models have been trained on datasets designed for 3D spatial understanding, such as absolute distance estimation and relative location inference. We benchmark SpatialRGPT and two checkpoints trained with the Spatial-VLM framework Chen et al. (2024): SpaceThinker-Qwen2.5-3B remyxai (2024) and SpaceQwen2.5-3B. The SpatialRGPT Cheng et al. (2024) model assumes that specific regions of the image are annotated with bounding boxes and is trained to answer 3D questions based on these regions. To evaluate SpatialRGPT on our Ego3D-Bench, we reformat the input: object names in the questions are replaced with placeholder labels (e.g., region-i), and a list of corresponding region captions is passed to the REC model to generate bounding boxes. These estimated bounding boxes were then overlaid on the images before being fed to SpatialRGPT.
@@
This category enhances the base VLM with a metric depth estimator and a REC model. We pass each query to the REC model Research (2023) to extract bounding boxes of the referred objects, and use a metric depth estimator Contributors (2024) to estimate depth values. We then construct a list that pairs each matched referring expression with its corresponding bounding box and estimated depth (i.e., distance from the ego). An example entry in this list is: [Front-View: Detected pedestrian with red hat at bbox [x1, y1, x2, y2], depth: z, Back-View:...] (examples in the Appendix). As shown in Table 4, equipping the VLM with REC and depth estimator models improves its baseline performance, highlighting the base model’s limitations in accurately identifying objects and estimating their depths. However, even with these enhancements, it falls short compared to our proposed Ego3D-VLM framework. This underscores the importance of integrating depth information into a unified map representation, enabling the VLM to reason more effectively about the 3D relationships between objects across different views.
Table 4: Ego3D-VLM vs. generalist VLMs empowered with REC and depth tools on Ego3D-Bench.
@@
Ego
Obj.
Ego
Obj.
Travel
Ego
Obj.
Ego
Obj.
Model
Dist.
Dist.
Loc.
Mot.
Mot.
Time
Rel.
Rel.
[…]
Dist.
Dist.
[…]
InternVL3-8B 25.8 28.7 29.8 54.1 54.8 36.1 65.2 49.9 43.1 15.2 39.1 27.2
InternVL3-8B+Depth+REC 60.9 46.9 35.2 58.8 49.0 31.9 69.1 60.7 51.6 6.6 19.6 13.1
Ego3D-VLM InternVL3-8B 65.3 56.1 36.9 69.0 74.0 47.5 67.4 63.0 60.1 6.8 9.0 8.0
Qwen2.5-7B 32.7 31.5 30.5 45.9 44.0 34.5 43.2 66.5 41.1 25.1 35.5 30.3
Qwen2.5-7B+Depth+REC 53.6 37.6 33.2 54.8 50.4 36.1 65.3 63.8 49.4 7.5 16.0 11.8
@@ 5.4 Ego3D-VLM on Further Benchmarks
Our proposed Ego3D-VLM is designed for ego-centric multi-view scenarios, which are particularly relevant for AI agent applications. In this section, however, we evaluate the performance of Ego3D-VLM in alternative multi-view settings, specifically those used in All-Angle Bench Yeh et al. (2025) and VSI-Bench Yang et al. (2024). Figure 7 illustrates the configurations of these benchmarks in comparison to Ego3D-Bench. All-Angle Bench features stationary multi-view cameras observing the same scene from different angles—a setup commonly used in motion capture and surveillance systems. VSI-Bench, on the other hand, involves a single moving camera within a static indoor environment, typically used for static scene reconstruction and spatial understanding. In contrast, our Ego3D-Bench assumes multiple cameras mounted on a moving ego agent, capturing the surrounding environment—a configuration suited for AI agent applications such as robotics and autonomous driving. Table 5 presents the performance of Ego3D-VLM compared to open- and closed-source baseline VLMs on All-Angle Bench and VSI-Bench. Despite the differences in data settings compared to our primary benchmark, Ego3D-VLM still outperforms the respective baselines, demonstrating its adaptability across diverse multi-view scenarios.
@@
InternVL3-38B 54.4 41.1
Ego3D-VLMInternVL3-38B 55.7 42.4
Table 6: Ablation on the main components of Ego3D-VLM. v4 includes all of the main components. v5 and v6 are shown for further analysis.
@@
5.5 Ablation Studies
Components of Ego3D-VLM. Table 6 presents an ablation study on the core components of Ego3D-VLM. Starting from the baseline, v0, we incrementally add each component to investigate its contribution. In version v1, we incorporate a cognitive map with estimated rotation ($R$), translation ($T$), and intrinsic parameters ($K$). Specifically, all cameras are positioned at the coordinate center with $T=[0,0,0]$. The front camera uses an identity rotation matrix, while all the other cameras—front-right, right, back-right, back, etc. —are incrementally rotated 45° around the Y-axis. The focal length is estimated from the images’ approximate field of view. In v2, we use the actual $K$ and in v3 we further use actual $R$, and $T$ relative to the front camera, i.e., often available as a fixed parameter for an embodied AI agent. v3 and v1 obtain a comparable RMSE, while v3 is only 2.4% higher in multi-choice QAs. Thus, even with estimated camera parameters our cognitive map can significantly boost the baseline’s performance. v4 adds relational scaling to cognitive map which decreases the RMSE by 2.5 meters. v4 is indeed Ego3D-VLM with all components. We provide two more ablations to evaluate the upper-bound of Ego3D-VLM. v5 assumes that the list of objects (only their names) in the input query are provided which enhances the REC results. v6 is the ground-truth cognitive map with ground-truth 3D locations.The difference between v6 and human level is only 5% showing the upper bound of Ego3D-VLM.
Table 7: Experiments on disentanglement of perception from reasoning
@@ 5.5 Ablation Studies
Disentanglement of Perception from Reasoning. In order to disentangle perception from reasoning, we have added the ground-truth 2D bounding boxes (BBox) for all objects, to the benchmark. This allows repeating our experiments with bounding boxes overlaid, effectively isolating the reasoning and perception components. As expected, adding GT BBox as visual prompts to the images will help the VLM to handle the perception procedure easier, therefore improving the results. The results in Table 7 show that the major limitation of baselines is not 2D perception, but 3D perception and reasoning.
Table 8: Ablation on the cognitive map format.
@@
Format of Cognitive Map. We explore three different formats for our generated cognitive map: visual, JSON, and textual. Examples of the visual and textual maps are shown in Figure 5. The average results for each format are presented in Table 8. Among the three, the textual and JSON formats achieve the best overall performance.
@@ 6 Conclusion and Future Work
In this work, we proposed Ego3D-Bench, a benchmark for spatial understanding of VLMs on ego-centric multi-view images. Overall, the benchmark shows a significant gap between human scores and VLMs. To address this limitation, we provided a post-training solution named Ego3D-VLM to enhance the performance of VLMs. Future work should explore fine-tuning VLMs with ego-centric multi-view QAs and incorporating 3D projection modules proposed in Ego3D-VLM in the course of fine-tuning. The limitations of this work will be provided in the Appendix.
@@ Appendix A Appendix
In this appendix, we present additional results on various source dataset splits of Ego3D-Bench, along with extended quantitative and qualitative analysis. We also report the performance of Ego3D-VLM on two public benchmarks, provide an inference time analysis, and discuss the limitations of our approach.
A.1 Ego3D-Bench: Further Details
@@
A.2 Results of different Source Datasets
NuScenes Waymo Argoverse
[…]
Closed-source Models
GPT-4o 60.9 21.2 60.0 19.0 59.4 17.1
Ego3D-VLM GPT-4o 73.5 8.5 73.1 7.5 72.7 6.2
Gemini-1.5-Pro 59.3 26.4 55.6 12.7 57.5 19.5
Ego3D-VLM Gemini-1.5-Pro 77.1 8.1 68.9 6.3 73.0 7.1
Qwen2.5 Family
Qwen2.5-3B 39.7 33.4 38.6 31.5 38.9 31.2
Ego3D-VLM Qwen2.5-3B 44.7 13.0 40.3 12.1 48.1 12.6
Qwen2.5-7B 40.9 36.0 40.6 24.7 41.6 30.1
Ego3D-VLM Qwen2.5-7B 56.4 11.3 51.9 8.3 54.4 8.7
Qwen2.5-32B 56.6 17.5 56.0 11.0 59.2 18.7
Ego3D-VLM Qwen2.5-32B 64.3 17.1 64.8 10.6 64.0 12.1
Qwen2.5-72B 58.2 21.2 57.7 11.1 58.2 16.2
Ego3D-VLM Qwen2.5-72B 74.4 8.5 64.4 6.4 69.5 7.6
InternVL3 Family
InternVL3-8B 42.3 28.3 44.0 21.7 42.9 31.5
Ego3D-VLM InternVL3-8B 62.0 10.5 57.1 6.5 61.3 6.7
InternVL3-14B 51.7 20.4 50.2 15.8 53.3 16.6
Ego3D-VLM InternVL3-14B 69.0 8.8 63.7 7.5 67.0 6.3
InternVL3-38B 51.2 31.4 50.5 19.4 53.4 25.4
Ego3D-VLM InternVL3-38B 71.7 8.7 64.3 8.2 67.0 6.3
InternVL3-78B 60.0 16.7 59.7 10.7 60.1 13.7
Ego3D-VLM InternVL3-78B 76.5 8.3 67.0 6.5 71.0 7.5
Ovis2 Family
Ovis2-4B 41.2 29.2 43.0 18.1 40.9 22.5
Ego3D-VLM Ovis2-4B 47.3 9.8 47.5 7.1 49.7 8.3
Ovis2-8B 41.5 21.4 41.9 20.7 41.0 20.2
Ego3D-VLM Ovis2-8B 55.6 9.4 55.1 6.8 57.9 7.0
Ovis2-16B 48.6 16.2 45.5 11.9 46.8 12.9
Ego3D-VLM Ovis2-16B 62.6 8.5 57.9 7.15 62.2 6.5
A.3 Benchmarking Generalist VLMs: Further Results
In the main paper, we reported results for InternVL3, Qwen2.5-VL, and Ovis-2 as representative open-source models, as they are, to the best of our knowledge, the current SOTA among open-source VLMs. Table 10 extends this comparison by including additional generalist VLMs on Ego3D-Bench, such as Gemini-1.5-Flash DeepMind (2023), Gemini-2-Flash DeepMind (2023), Gemini-2.5-Flash DeepMind (2023), Phi-3.5 et al. (2024), LLaVA-One-Vision-7B Li et al. (2024a), LLaVA-Next-Video-7B Zhang et al. (2024b), and DeepSeek-VL2 Wu et al. (2024c). We observe that LLaVA-One-Vision, Phi-3.5, and DeepSeek-VL2 underperform compared to InternVL3 and Qwen2.5-VL models of similar size. For instance, InternVL3-8B achieves an accuracy of 43%, while LLaVA-One-Vision-7B achieves 38.7%. Likewise, Phi-3.5 (3.8B) attains an accuracy of 40.3%, compared to 41% from Ovis-2 (4B).

### v2

@@ 5 Evaluation on Ego3D-Bench
In this section, we present a comprehensive evaluation of VLMs on Ego3D-Bench. We organize our experiments into 4 parts by benchmarking (1) generalist VLMs (i.e., models trained for general vision-language tasks), (2) 3D-VLMs (i.e., models trained specifically for 3D SU), (3) VLM+Depth+REC (i.e., generalist VLMs augmented with depth and REC tools), and (4) ablation studies. We use a fixed R1-style prompt [16] in the evaluations of all models (details in the Appendix A.8). We use Grounding-Dino-Base [35] as the REC model and Depth-Anything-V2-Metric-Large [7] as the metric depth estimator in all experiments.
Table 1: Comparison results of generalist VLMs vs. Ego3D-VLM on Ego3D-Bench.
@@
5.1 Benchmarking Generalist VLMs
We use GPT-4o [30] and Gemini-2-Flash [9] closed-source VLMs, and 3 competitive families of open-source models: InternVL3 [11], Qwen2.5-VL [32], and Ovis2 [26] open-source models. Inference time analysis, numerical results on more generalist VLMs, and qualitative examples are given in the Appendix A.5, A.4, and A.8.
@@ 5.1 Benchmarking Generalist VLMs
Figure 6: Average performance of leading VLMs w/ and w/o Ego3D-VLM vs. chance & human levels on each category of Ego3D-Bench.
Performance Analysis. Figure 6 shows the average performance of 4 leading models (GPT-4o, Gemini-1.5-Pro, InternVL3-78B, and Qwen2.5-VL-72B) with and without Ego3D-VLM integration across all multiple-choice QAs. Travel time, localization, and object-centric absolute distance are the most challenging tasks for VLMs with 40%-45% average accuracy. We argue that these categories require intricate spatial reasoning and the ability to construct an accurate mental map by the relative positioning of objects. While humans can simply build such maps and achieve perfect accuracy, VLMs struggle to replicate this level of SU, indicating a key area for further development. In the absolute distance tasks, the incorporation of 3D location data through the cognitive map substantially narrows the gap between model and human level accuracy. Notably, for object-centric absolute distance, VLMs augmented with Ego3D-VLM even surpass human. This is expected, as human estimation of 3D distances in object-centric cases is prone to substantial error without explicit 3D spatial info. Conversely, in the localization task, VLMs continue to fall short of human proficiency, even with cognitive map support.
Table 2: Blind Ego3D-VLM performance analysis.
@@
Blind VLM Performance. This baseline evaluates how much spatial reasoning can be achieved by VLMs using only textual input, without any visual information, relying solely on their world knowledge [29]. We report the average performance of GPT-4o and Gemini-1.5-Pro models. The blind VLMs perform 5% worse than vision-enabled VLMs (53.8% vs. 58.8%) and 16.4% better than chance level. These results are consistent with findings from prior single-view and video-based spatial reasoning benchmarks [6, 44].
Blind Ego3D-VLM Performance. we use the same models for fair comparison. Table 2 shows that Ego3D-VLM with VLM outperforms Ego3D-VLM with LLM since VLMs can ignore false positives in the cognitive map and remain robust when false negatives occur, whereas LLMs suffer performance drops in those cases.
@@ 5.2 Benchmarking 3D-VLMs
3D-VLM models have been trained on datasets designed for 3D SU, such as absolute distance estimation and relative location inference. We benchmark SpatialRGPT and two checkpoints trained with the Spatial-VLM framework [5]: SpaceThinker-Qwen2.5-3B [34] and SpaceQwen2.5-3B. The SpatialRGPT [6] model assumes that specific regions of the image are annotated with bounding boxes and is trained to answer 3D questions based on these regions. To evaluate SpatialRGPT on our Ego3D-Bench, we reformat the input: object names in the questions are replaced with placeholder labels (e.g., region-i), and a list of corresponding region captions is passed to the REC model to generate bounding boxes. These estimated bounding boxes were then overlaid on the images before being fed to SpatialRGPT.
@@
Table 4: Ego3D-VLM vs. generalist VLMs with REC+depth tools on Ego3D-Bench.
Mult. Choice Abs. Ans.
@@
InternVL3-8B 43.1 27.2
InternVL3-8B+Depth+REC 51.6 13.1
Ego3D-VLM InternVL3-8B 60.1 8.0
Qwen2.5-7B 41.1 30.3
Qwen2.5-7B+Depth+REC 49.4 11.8
Ego3D-VLM Qwen2.5-7B 54.3 9.5
This category enhances the base VLM with a metric depth estimator and a REC model. We pass each query to the REC model [35] to extract BBox of the referred objects, and use a metric depth estimator [7] to estimate depth values. We then construct a list that pairs each matched referring expression with its corresponding BBox and estimated depth (i.e., distance from the ego). An example entry in this list is: [Front-View: Detected pedestrian with red hat at bbox [x1, y1, x2, y2], depth: z, Back-View:...] (examples in the Appendix A.8). As shown in Table 4, equipping the VLM with REC and depth estimator models improves its baseline performance, highlighting the base model’s limitations in accurately identifying objects and estimating their depths. However, even with these enhancements, it falls short compared to our proposed Ego3D-VLM framework. This underscores the importance of integrating depth information into a unified map representation, enabling the VLM to reason more effectively about the 3D relationships between objects across different views.
@@ 5.4 Ego3D-VLM on Further Benchmarks
Our Ego3D-VLM is designed for ego-centric multi-view cases, which are particularly relevant for AI agent applications. In this section, however, we evaluate the performance of Ego3D-VLM in alternative multi-view settings, specifically those used in All-Angle Bench [46] and VSI-Bench [44]. Table 5 presents the performance of Ego3D-VLM compared to open- and closed-source baseline VLMs on All-Angle Bench and VSI-Bench. Despite the differences in data settings compared to our primary benchmark, Ego3D-VLM still outperforms the respective baselines, demonstrating its adaptability across diverse multi-view scenarios. More details are given in the Appendix A.5.
@@
Table 6: Ablation on Ego3D-VLM main components. v4: all components.
@@
Table 7: Perception-reasoning disentanglement.
@@
5.5 Ablation Studies
Ego3D-VLM Components. Table 6 presents an ablation on Ego3D-VLM core components. Starting from the baseline, v0, we incrementally add each component. In v1, we use a cognitive map with estimated rotation ($R$), translation ($T$), and intrinsic params ($K$). Specifically, all cameras are positioned at the coordinate center with $T=[0,0,0]$. The front camera uses an identity rotation matrix, while all the other cameras—front-right, right, back-right, back, etc. —are incrementally rotated 45° around the Y-axis. The focal length is estimated from the images’ approximate field of view. In v2, we use the actual $K$ and in v3 we further use actual $R$, and $T$ relative to the front camera, i.e., often available as a fixed parameter for an embodied AI agent. v3 and v1 obtain a comparable RMSE, while v3 is only 2.4% higher in multi-choice QAs. Thus, even with estimated camera parameters our cognitive map can significantly boost the baseline’s performance. v4 adds relational scaling to cognitive map which decreases the RMSE by 2.5 meters. v4 is indeed Ego3D-VLM with all components. We provide two more ablations to evaluate the upper-bound of Ego3D-VLM. v5 assumes that the list of objects (only their names) in the input query are provided which enhances the REC results. v6 is the ground-truth cognitive map with ground-truth 3D locations.The difference between v6 and human level is only 5% showing the upper bound of Ego3D-VLM.
Table 8: Ablation on cognitive map format.
@@ 5.5 Ablation Studies
Perception-Reasoning Disentanglement. In order to disentangle perception from reasoning, we add the ground-truth 2D BBox for all objects to the benchmark. This allows repeating our experiments with BBox overlaid, effectively isolating the reasoning and perception. As expected, adding GT BBox as visual prompts to the images will help the VLM to handle perception easier, therefore improving the results. The results in Table 7 show that the major limitation of baselines is not 2D perception, but 3D perception and reasoning. More reasoning analysis on Ego3D-Bench as well as Ego3D-VLM’s robustness in challenging conditions are given in Appendix A.3 and A.6.
Cognitive Map Format. We explore 3 different formats for our generated cognitive map: visual, JSON, and textual. Examples of the visual and textual maps are shown in Figure 5. The average results are presented in Table 8. Among the three, the textual and JSON formats achieve the best overall performance.
@@ 6 Conclusion and Future Work
In this work, we proposed Ego3D-Bench, a benchmark for spatial understanding of VLMs on ego-centric multi-view images. Overall, the benchmark shows a significant gap between human scores and VLMs. To address this limitation, we provided a post-training solution named Ego3D-VLM to enhance the performance of VLMs. Future work should explore fine-tuning VLMs with ego-centric multi-view QAs and incorporating 3D projection modules proposed in Ego3D-VLM in the course of fine-tuning. The limitations of this work are provided in the Appendix (A.9).
@@ Appendix A Appendix
In this appendix, we present the following additional discussions and experimental results corresponding to our proposed Ego3D-Bench and Ego3D-VLM:
  • Ego3D-Bench vs. Other Benchmarks
  • Further details on Ego3D-Bench creation and question templates
  • Benchmarking the thinking of Ego3D-Bench as an Open-QA
  • Results of the three source datasets: NuScenes, Waymo, and Argoverse
  • Benchmarking more generalist VLMs
  • Inference time analysis of Ego3D-VLM vs. other baselines
  • Robustness of Ego3D-VLM in challenging conditions
  • Qualitative results and prompts structure
  • Limitations and future work
A.1 Ego3D-Bench vs. Other Benchmarks
As described in the main body of this paper, our proposed Ego3D-VLM is designed for ego-centric multi-view scenarios, which are particularly relevant for AI agent applications. Ego3D-VLM is different from alternative multi-view settings, specifically those used in All-Angle Bench [46] and VSI-Bench [44]. Figure 7 illustrates the configurations of these benchmarks in comparison to Ego3D-Bench. All-Angle Bench features stationary multi-view cameras observing the same scene from different angles—a setup commonly used in motion capture and surveillance systems. VSI-Bench, on the other hand, involves a single moving camera within a static indoor environment, typically used for static scene reconstruction and SU. In contrast, our Ego3D-Bench assumes multiple cameras mounted on a moving ego agent, capturing the surrounding environment—a configuration suited for AI agent applications such as robotics and autonomous driving.
A.2 Ego3D-Bench: Further Details
@@
A.3 Benchmarking the Thinking of Ego3D-Bench as an Open-QA
Thinking of Obj. Abs. Dist. Thinking of Ego Abs. Dist.
$\uparrow GPT-4o Score (0-10)\uparrow$
InternVL3-8B 1.9 2.0
Ego3D-VLM InternVL3-8B 4.7 5.2
InternVL3-14B 2.3 2.4
Ego3D-VLM InternVL3-14B 5.7 5.0
InternVL3-38B 2.1 2.5
Ego3D-VLM InternVL3-38B 7.1 3.7
InternVL3-78B 2.1 1.7
Ego3D-VLM InternVL3-78B 7.0 6.8
A.4 Benchmarking Generalist VLMs: Further Results
In the main paper, we reported results for InternVL3, Qwen2.5-VL, and Ovis-2 as representative open-source models, as they are, to the best of our knowledge, the current SOTA among open-source VLMs. Table 10 extends this comparison by including additional generalist VLMs on Ego3D-Bench, such as Gemini-1.5-Flash [9], Gemini-2-Flash [9], Gemini-2.5-Flash [9], Phi-3.5 [12], LLaVA-One-Vision-7B [21], LLaVA-Next-Video-7B [48], and DeepSeek-VL2 [43]. We observe that LLaVA-One-Vision, Phi-3.5, and DeepSeek-VL2 underperform compared to InternVL3 and Qwen2.5-VL models of similar size. For instance, InternVL3-8B achieves an accuracy of 43%, while LLaVA-One-Vision-7B achieves 38.7%. Likewise, Phi-3.5 (3.8B) attains an accuracy of 40.3%, compared to 41% from Ovis-2 (4B).